# Nonlinear Model Order Reduction of Dynamical Systems in Process Engineering: Review and Comparison

### v1

@@ 6.  Discussion
Based on theoretical properties of the MOR methods and the findings of the case study, we summarize characteristic properties of the methods investigated. However, we emphasize that definitive statements on a method ranking are delicate, because in addition to $n_{z}$, most MOR techniques possess further hyperparameters that can strongly affect ROM behavior, e.g., the position of the sensitive element (COMP, AGG) or the ANN architecture (MFL, KW, MFL-ANN). Although we performed a basic validation of crucial hyperparameters within the case study, we do not claim that this screening was exhaustive. Additionally, the performance of a certain type of ROM may be system-dependent, and in that sense our experimental results may be specific to air separation.
@@
Comparing the results of intrusive (MFL, MFLu) to non-intrusive nonlinear subspace methods (KW, MFL-ANN) in the case study, we conclude that similar ROM accuracies are reached when targeting very low-order models. Such very-low-order ROMs are important in control applications, where CPU reduction has a high priority whereas moderate errors are acceptable. Often, a moderate prediction error can be compensated by state estimation and model adaptation, especially if the error is predominantly steady-state offset [243]. On the other hand, despite a relatively low RMSE, the POD-Res results exhibited an IR in the very-low-order setup, which can be problematic in control applications.
@@
When prioritizing accuracy over low order, non-intrusive methods may struggle to reach arbitrary precision due to their inherently difficult learning problems. In such cases, intrusive strategies can be more easy to apply. In particular, the POD-based methods were able to generate very accurate ROMs given a sufficiently high order. In this context, we acknowledge the simplicity and approximation guarantees (on the training data) of the POD. In contrast, high projection and prediction accuracies were difficult to reach when using MFL strategies.
@@
The advantage of non-intrusive over intrusive reduction methods is that non-intrusive strategies combine MOR and hyperreduction. The weakness of non-intrusive methods is that they can be prone to overfitting if not carefully regularized. One possible “structural regularization” strategy is to restrict the latent dynamics to a linear or bilinear form. In particular, applied Koopman theory combines MFL and linearization into a KW structure, which enables reliable ROMs of the moderately nonlinear systems.
@@
POD-Res and COMP may be considered as data-driven and heuristic variants of SPT, respectively. However, COMP retains most of the model sparsity, whereas POD-Res ROMs are generally dense and can therefore be more challenging to solve. Here, both POD-Res and COMP were numerically the most delicate to handle. Specifically, large-scale implicit nonlinear systems of equations and non-phyiscal IR made these models difficult to converge.
@@
In terms of reduction effort, the (non-intrusive) KW and MFL-ANN strategies involved the most extensive offline computations. Achieving highly accurate models may require multiple trainings and extensive hyperparameter tuning, which introduces delays into the model reduction procedure. Furthermore, non-intrusive methods are most sensible to the training data in our experience. In particular, a balanced amount of transient and stationary data is essential for the successful generation of ROMs that enable precise long-term forecasts.
@@
The model reduction efforts in intrusive reduction methods (including tailored methods) demand a great share of model building and reformulation (debugging, robustification). On the other hand, the ML problems used within intrusive methods, e.g., AEN training, are typically less computationally demanding than in non-intrusive approaches. In our experience, intrusive data-driven methods are slightly more tolerant to a poorly designed data set, e.g., too few steady-state data. For intrusive methods, further efforts have to be expected in terms of hyperreduction, often necessary to obtain a real-time capable ROM.
@@
Lastly, we comment on the model reduction procedure for interconnected systems. Applying the tailored methods, such as COMP and AGG, is naturally done in a “subsystem-by-subsystem” fashion, since these methods provide reduced models of the subsystems, e.g., distillation columns or HXs, of an interconnected system, e.g., a process flowsheet model. In particular, when working in a structured modeling environment, such as Unisim Design or gPROMS, these ROMs can be pre-configured and included into the model library. Reducing an interconnected system model is then achieved by substituting full-order submodels by reduced submodels. The “subsystem-by-subsystem” approach preserves the model sparsity originating from the interconnection, but the feasible degree of reduction may be limited, as shown by the case study results. Further, the reduction problem becomes a combinatorial problem due to potentially many subsystem ROMs to be configured and combined.
@@
In contrast to the “subsystem-by-subsystem” approach, general-purpose methods are most easily applied to the entire FOM. Then, the complete model of the interconnected system is reduced as a whole. As shown by the case study, such an approach can enable a lower ROM order. However, for systems with many inputs, the data sampling efforts grow exponentially with $n_{u}$, which can result in a curse of dimensionality. This curse may be broken by a “subsystem-by-subsystem” strategy if the high-dimensional inputs enter the system only locally and are evenly distributed over the submodels.
@@ 7.  Conclusions and outlook
Model order reduction spans a collection of techniques to construct a low-order model approximating the evolution of a high-order system. Herein, we have reviewed and compared reduction methods for nonlinear dynamical models with inputs. In particular, we have contrasted the different strategies both theoretically and in a case study. As expected, there is not universal best choice when selecting a reduction method. Rather, a suitable choice depends on targeted ROM properties.
@@
If very-low-order ROMs are needed, then intrusive nonlinear subspace methods, e.g., the MFLu approach, as well as non-intrusive nonlinear subspace methods, e.g., the KW model, have provided accurate ROMs within our case study. Notably, non-intrusive reduction methods combine MOR and hyperreduction in a single step, providing a ROM that is real-time capable. If a high ROM accuracy is needed but the degree of reduction is of secondary importance, then linear subspace methods, e.g., the POD-Galerkin method, are a good choice due to their (theoretically) arbitrary degree of accuracy and simplicity of application. Finally, when desiring structural preservation of a given (network or flowsheet) model comprising interconnected subsystems, then problem-specific ROMs are a straightforward choice to reduce the subunits individually by means of pre-configured ROM blocks.
@@
Besides reviewing and comparing existing MOR methods, we have also proposed an extension of the manifold-Galerkin method to systems with inputs (Section 3.1.5). In particular, we stress that system inputs $\bm{u}(t)$ can deform the slow manifold $\mathcal{M}_{\mathcal{X}}$. Consequently, if the effect of inputs is disregarded, an accurate approximation of $\mathcal{M}_{\mathcal{X}}$ may necessitate a higher ROM order. Within the case study, we validated this extension and found that our approach outperforms the state-of-the-art manifold-Galerkin method in cases of very strong reduction both in terms of the projection accuracy and the ROM prediction accuracy. As expected, our extended approach and the established approach performed similarly at higher orders.
@@
In our case study, we did not investigate hyperreduction or CPU cost of integration. Instead, we expect that a sufficiently low order combined with hyperreduction enables real-time applications, e.g., [183, 120]. Clearly, real-time capability depends on the application, so there is no universal CPU threshold to classify performance. In any case, accounting for hyperreduction already within MOR could be a promising step towards more holistic methods.
@@
Despite increasing computer power, nonlinear MOR (especially for parametric and non-autonomous systems) remains a challenging discipline worth of investigation. The curse of dimensionality, when sampling systems with many inputs or parameters, is connected to the need for robust methods applicable in the low-data limit. In this regard, methods such as physics-informed strategies [244, 104] are a promising approach to enrich the non-intrusive manifold learning task by additional (possibly intrusive) information. Moreover, applying manifold learning methods to very-high-order problems remains an open challenge.
Unfortunately, there are no general guarantees for the preservation of stability properties by MOR methods [34]. Consequently, a-posteriori error and stability analysis are currently required when conducting model reduction. Hence, methods with a-priori error estimates or even guarantees are needed. Despite some recent advances towards modal closure and memory effects, e.g., using the Mori-Zwanzig formalism [245], these works are at an early stage [246].
@@
Finally, we notice some limitations of current reduction methods with respect to numerical optimization and control applications. First, most numerical optimization techniques are based on derivative information, in particular, the Jacobian of the ROM variables $\bm{z}(t)$ and $\bm{x}_{\mathrm{r}}(t)$ in Eq. 4 with respect to each other and the inputs $\bm{u}(t)$. However, standard MOR methods do not explicitly account for the accuracy of such derivatives. For learning-based MOR and hyperreduction approaches, Sobolev trainings [247, 248] may improve derivative accuracy. Additionally, in our experience, sensitivity integration is often considerably more expensive than standard integration, and therefore accounting for sensitivities in the reduction process might enable fundamental improvements in CPU costs of optimization applications. Second, model reduction of non-smooth or even discontinuous models is interesting, e.g., for controlling a process start-up, but remains a mostly unexplored topic [249, 250].
@@ Acknowledgments
We thank Danimir Doncevic and Eike Cramer for fruitful discussions.

### v2

@@ 6. Discussion
Based on theoretical properties of the MOR methods and the findings of the case study, we summarize characteristic properties of the methods investigated. However, we emphasize that definitive statements on method ranking are delicate, because in addition to $n_{z}$, most MOR techniques possess further hyperparameters that can strongly affect ROM behavior, e.g., the position of the sensitive element (COMP, AGG) or the ANN architecture (MFL, KW, MFL-ANN). Although we performed a basic validation of crucial hyperparameters within the case study, we do not claim that this screening was exhaustive. Additionally, the performance of a certain type of ROM may be system-dependent, and in that sense our experimental results may be specific to air separation units.
@@
Comparing the results of intrusive (MFL, MFLu) to non-intrusive nonlinear subspace methods (KW, MFL-ANN) in the case study, we conclude that similar ROM accuracy is reached when targeting very low-order models. Such very-low-order ROMs are important in control applications, where CPU reduction has a high priority whereas moderate errors are acceptable. Often, a moderate prediction error can be compensated by state estimation and model adaptation, especially if the error is predominantly steady-state offset [196]. On the other hand, despite a relatively low RMSE, the POD-Res results exhibited an inverse response in the very-low-order setup, which can be problematic in control applications.
@@
When prioritizing accuracy over low order, non-intrusive methods may face difficulties to reach arbitrary precision due to their inherently difficult learning problems. In such cases, intrusive strategies can be easier to apply. In particular, the POD-based methods were able to generate very accurate ROMs given sufficiently high order. In this context, we acknowledge the simplicity and approximation guarantees (on the training data) of the POD. In contrast, high projection and prediction accuracies were difficult to reach when using MFL strategies.
@@
The advantage of non-intrusive over intrusive reduction methods is that non-intrusive strategies combine MOR and hyperreduction. The weakness of non-intrusive methods is that they can be prone to overfitting if not carefully regularized. One possible structural regularization strategy is to restrict the latent dynamics to a linear or bilinear form. In particular, applied Koopman theory combines MFL and linearization into a KW structure, which enables reliable ROMs of the moderately nonlinear systems.
@@
POD-Res and COMP may be considered as data-driven and heuristic variants of SPT, respectively. However, COMP retains most of the model sparsity, whereas POD-Res ROMs are generally dense and can therefore be more challenging to solve. Here, both POD-Res and COMP were numerically the most delicate to handle. Specifically, large-scale implicit nonlinear systems of equations and non-physical inverse response made these models difficult to converge.
@@
In terms of reduction effort, the (non-intrusive) KW and MFL-ANN strategies involved the most extensive offline computations. Achieving sufficiently accurate models may require multiple trainings and extensive hyperparameter tuning, which introduces delays into the model reduction procedure. Furthermore, non-intrusive methods are most sensible to the training data in our experience. In particular, a balanced amount of transient and stationary data is essential for the successful generation of ROMs that enable precise long-term forecasts.
@@
The model reduction efforts in intrusive reduction methods (including tailored methods) demand a great share of model building and reformulation (debugging, robustification). On the other hand, the training problems used within intrusive methods, e.g., AEN training, are typically less computationally demanding than in non-intrusive approaches. In our experience, intrusive data-driven methods are slightly more tolerant to a poorly designed data set (e.g., too few steady-state data) than non-intrusive methods. For intrusive methods, further efforts have to be expected in terms of hyperreduction, often necessary to obtain a real-time capable ROM.
@@
Lastly, we comment on the model reduction procedure for interconnected systems. Applying the tailored methods, such as COMP and AGG, is naturally done in a subsystem-by-subsystem fashion, since these methods provide reduced models of the subsystems, e.g., distillation columns or HXs, of an interconnected system, e.g., a process flowsheet model. In particular, when working in a structured modeling environment, such as Unisim Design or gPROMS, these ROMs can be pre-configured and included into the model library. Reducing an interconnected system model is then achieved by substituting full-order submodels by reduced submodels. The “subsystem-by-subsystem” approach preserves the model sparsity originating from the interconnection. However, the feasible degree of reduction may be limited, as shown by the case study results. Further, the reduction problem becomes a combinatorial problem due to potentially many subsystem ROMs to be configured and combined.
From a mathematical perspective, the subsystem-by-subsystem strategy falls into the port-Hamiltonian formalism. Notably, in this framework, the respective ROMs are invariant under couplings and preserve their properties. Hence, this formalism should be investigated in future work to develop more accurate alternatives to COMP and AGG.
@@
In contrast to the subsystem-by-subsystem approach, general-purpose methods are most easily applied to the entire FOM. Then, the complete model of the interconnected system is reduced as a whole. As shown by the case study, such an approach can enable a lower ROM order. However, for systems with many inputs, the data sampling efforts grow exponentially with $n_{u}$, which can result in a curse of dimensionality. This curse may be broken by a subsystem-by-subsystem strategy if the high-dimensional inputs enter the system only locally and are evenly distributed over the submodels.
@@ 7. Conclusions and outlook
Model order reduction spans a collection of techniques to construct a low-order model approximating the evolution of a high-order system. Herein, we have reviewed and compared reduction methods for nonlinear dynamical models with inputs. In particular, we have contrasted the different strategies both theoretically and in a case study. As expected, there is no universal best choice when selecting a reduction method. Rather, a suitable choice depends on targeted ROM properties.
@@
If very-low-order ROMs are needed, then intrusive nonlinear subspace methods, e.g., the MFLu approach, as well as non-intrusive nonlinear subspace methods, e.g., the KW model, have provided accurate ROMs within our case study. Notably, non-intrusive reduction methods combine MOR and hyperreduction in a single step, providing a ROM that is real-time capable. If high ROM accuracy is needed but the degree of reduction is of secondary importance, then linear subspace methods, e.g., the POD-Galerkin method, are a good choice due to their (theoretically) arbitrary degree of accuracy and simplicity of application. Finally, when desiring structural preservation of a given network or flowsheet model comprising interconnected subsystems, then structure-preserving (especially port-Hamiltonian) ROMs or problem-specific ROMs are a straightforward choice. These approaches reduce the subunits individually by means of pre-configured ROM blocks.
@@
Besides reviewing and comparing existing MOR methods, we have also proposed an extension of the manifold-Galerkin method to systems with inputs (Section 3.1.5). In particular, we stress that system inputs $\bm{u}(t)$ can deform the slow manifold $\mathcal{M}_{\mathcal{X}}$. Consequently, if the effect of inputs is disregarded, an accurate approximation of $\mathcal{M}_{\mathcal{X}}$ may necessitate a high ROM order.
Within the case study, we validated the proposed manifold-Galerkin extension for systems with inputs. In cases of very strong reduction, we found that our approach outperforms the standard method both in terms of the projection accuracy and the ROM prediction accuracy. As expected, our extended approach and the established approach performed similarly at higher orders.
@@
In our case study, we did not investigate hyperreduction or CPU cost of integration. Instead, we expect that a sufficiently low order combined with hyperreduction enables real-time applications, e.g., [254, 260]. Clearly, real-time capability depends on the application, so there is no universal CPU threshold to classify performance. In any case, accounting for hyperreduction already within MOR could be a promising step towards more holistic methods. Even though our case study represents a realistic MOR problem in the process systems control field, future work should investigate higher-dimensional problems and examine additional reduction methods, e.g., [39, 77].
@@
Despite increasing computer power, nonlinear MOR (especially for parametric and non-autonomous systems) remains a challenging discipline worth of investigation. The curse of dimensionality, when sampling systems with many inputs or parameters, is connected to the need for robust methods applicable in the low-data limit. In this regard, structure-preserving MOR is a promising approach to enrich the non-intrusive manifold learning task by additional (physics-based) information. Moreover, the scalability of manifold learning methods to very-high-order problems remains an open challenge. Recently, promising results have been presented using quadratic/polynomial manifold learning [279, 77] and convolutional autoencoders [74, 246].
Although several methods offer stability preservation, such guarantees are not inherently provided by every method [117]. Consequently, a careful method selection or a-posteriori error and stability analysis are required. Another important topic is the development of a-priori error bounds on ROM accuracy.
@@
Finally, we notice some limitations of current reduction methods with respect to numerical optimization and control applications. First, most numerical optimization techniques are based on derivative information, in particular, the Jacobian of the ROM variables $\bm{z}(t)$ and $\bm{x}_{\mathrm{r}}(t)$ in (4) with respect to each other and the inputs $\bm{u}(t)$. However, standard MOR methods do not explicitly account for the accuracy of such derivatives. For learning-based MOR and hyperreduction approaches, Sobolev trainings [54, 282] may improve derivative accuracy. Additionally, in our experience, sensitivity integration is often considerably more expensive than standard integration, and therefore accounting for sensitivities in the reduction process might enable fundamental improvements in CPU costs of optimization applications. Second, model reduction of non-smooth or even discontinuous models, e.g., for controlling a process start-up, is a relevant but mostly unexplored topic [304, 27].
@@ Acknowledgments
A pre-print version of this manuscript has been posted on https://arxiv.org/pdf/2506.12819 under the arXiv.org non-exclusive distribution license. The authors gratefully acknowledge the financial support of the Kopernikus project SynErgie by the German Federal Ministry of Research, Technology and Space (BMFTR) and project supervision by Projektträger Jülich (PtJ). The authors thank Danimir Doncevic and Eike Cramer for fruitful discussions. Finally, the authors thank the anonymous reviewers and Tony Roberts for their valuable feedback and suggestions.